OpenStudy (anonymous):

linear approximation of the volume of a cone - V=(1/3)pi x r^2 x h

OpenStudy (ahsome):

Rewritten Equation: \[V=\frac{1}{3}\pi^2h\]

OpenStudy (ahsome):

What is the question?

OpenStudy (anonymous):

@Ahsome

OpenStudy (anonymous):

i have found that r=6 and h=6

OpenStudy (anonymous):

this z is surface area??

OpenStudy (anonymous):

i thought z=h?? I'm probably wrong though @aryandecoolest

OpenStudy (anonymous):

nopes surely z is not equal to h..... this function describes it's surface area i guess..

OpenStudy (phi):

to figure out the radius I would write the equation as \[ \sqrt{4x^2 +8x + 4y^2 -24y +40}= 6-z \] or after factoring a 4 from the square root (making it a 2) \[ \sqrt{x^2 +2x + y^2 -6y +10}= \frac{6-z}{2} \\ x^2 +2x + y^2 -6y +10 = \frac{(6-z)^2}{4} \] complete the square on both the x and y, by writing 10 as 1+9 and putting 1 and 9 in the "correct spots": \[ x^2 +2x +1 + y^2 -6y +9 = \frac{(6-z)^2}{4} \\ (x+1)^2 + (y-3)^2 = \frac{(6-z)^2}{4} \] that is the equation of a circle with center (-1,3) and radius \[ r = \frac{6-z}{2} \] the radius when z =0 (i.e. on the ground), is 6/2= 3

OpenStudy (anonymous):

thanks @phi what would h be?

OpenStudy (phi):

your "h" is correct, h=6

OpenStudy (anonymous):

ah right, and r=3, thanks! @phi